# A relativistic model of b-EMRI systems and their gravitational radiation

### v1

@@ Appendix B Teukolsky source term
The source term in Eq. (12) is given by $\hat{T}=2\rho^{-4}T_{4}$, where $T_{4}$ a combination of the projected stress-energy tensor defined by
@@ Appendix C Analytical expressions of split amplitudes
$Z^{(1)}_{lm}(\tau,\omega)=$$\big{\{}\tilde{\alpha}_{1}\sin\left[\left(\Omega+\omega_{IB}\right)\tau+\tilde{\psi}\right]$ (82)
$+\tilde{\alpha}_{2}\cos\left[\left(\Omega+\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+\tilde{\alpha}_{3}\sin\left[\left(\Omega-\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+\tilde{\alpha}_{4}\cos\left[\left(\Omega-\omega_{IB}\right)\tau+\tilde{\psi}\right]$
$+\tilde{\alpha}_{5}\sin\left(\omega_{IB}\tau\right)+\tilde{\alpha}_{6}\cos\left(\omega_{IB}\tau\right)\big{\}}$
@@
$Z^{(1)}_{lm}(\tau,\omega)=$ $Z^{(1)}_{lm-1-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\Omega+\omega_{IB}\right)\tau\right]$ (83)
$+Z^{(1)}_{lm11}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\Omega-\omega_{IB}\right)\tau\right]$
$+Z^{(1)}_{lm-11}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\Omega-\omega_{IB}\right)\tau\right]$
$+Z^{(1)}_{lm1-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\Omega+\omega_{IB}\right)\tau\right]$
$+Z^{(1)}_{lm0-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\omega_{IB}\right)\tau\right]$
$+Z^{(1)}_{lm01}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\omega_{IB}\right)\tau\right].$
@@
$Z^{(1)}_{lm-1-1}=\frac{-i\tilde{\alpha}_{1}+\tilde{\alpha}_{2}}{2}e^{i\tilde{\psi}},$ (84a)
$Z^{(1)}_{lm-11}=\frac{-i\tilde{\alpha}_{3}+\tilde{\alpha}_{4}}{2}e^{i\tilde{\psi}},$ (84b)
@@
$Z^{(1)}_{lm1-1}=\frac{i\tilde{\alpha}_{3}+\tilde{\alpha}_{4}}{2}e^{-i\tilde{\psi}},$ (84e)
$Z^{(1)}_{lm11}=\frac{i\tilde{\alpha}_{1}+\tilde{\alpha}_{2}}{2}e^{-i\tilde{\psi}},$ (84f)
@@
$\tilde{\alpha}_{1}=$ $\frac{\gamma\left(1-\sin\tilde{\theta}\right)}{4\gamma_{\varphi\varphi}^{1/2}}\left[\alpha_{1}\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{\mathcal{M}}+\alpha_{4}\right]$
$-\frac{(\Omega+\omega_{IB})(1-\sin\tilde{\theta})}{4\gamma_{rr}^{1/2}}\alpha_{5},$ (85a)
$\tilde{\alpha}_{2}=$ $\frac{1-\sin\tilde{\theta}}{4\gamma_{rr}^{1/2}}\alpha_{2},$ (85b)
$\tilde{\alpha}_{3}=$ $\frac{\gamma\left(1+\sin\tilde{\theta}\right)}{4\gamma_{\varphi\varphi}^{1/2}}\left[\alpha_{1}\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{\mathcal{M}}+\alpha_{4}\right]$
$-\frac{(\Omega-\omega_{IB})(1+\sin\tilde{\theta})}{4\gamma_{rr}^{1/2}}\alpha_{5},$ (85c)
$\tilde{\alpha}_{4}=$ $\frac{1+\sin\tilde{\theta}}{4\gamma_{rr}^{1/2}}\alpha_{2},$ (85d)
$\tilde{\alpha}_{5}=$ $\frac{\cos\tilde{\theta}}{2\gamma_{\theta\theta}^{1/2}}\alpha_{3},$ (85e)
$\tilde{\alpha}_{6}=$ $\frac{\omega_{IB}\cos\tilde{\theta}}{2\gamma_{\theta\theta}^{1/2}}\alpha_{6}.$ (85f)
@@
$c_{01}=$ $\vartheta_{11}\sin\left[(\Omega+\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{12}\sin\left[(\Omega-\omega_{IB})\tau+\tilde{\psi}\right],$ (87a)
$c_{11}=$ $\vartheta_{21}\cos\left[(\Omega+\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{22}\cos\left[(\Omega-\omega_{IB})\tau+\tilde{\psi}\right],$ (87b)
$c_{11}^{\prime}=$ $\vartheta_{21}^{\prime}\sin\left[(\Omega+\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{22}^{\prime}\sin\left[(\Omega-\omega_{IB})\tau+\tilde{\psi}\right],$ (87c)
$c_{21}=$ $\vartheta_{31}\sin(\omega_{IB}\tau),$ (87d)
$c_{21}^{\prime}=$ $\vartheta_{31}^{\prime}\cos(\omega_{IB}\tau),$ (87e)
$c_{31}=$ $\vartheta_{41}\sin\left[(\Omega+\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{42}\sin\left[(\Omega-\omega_{IB})\tau+\tilde{\psi}\right],$ (87f)
$c_{02}=$ $\vartheta_{51}\sin\left[2(\Omega+\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{52}\sin\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\vartheta_{53}\sin\left[2(\Omega-\omega_{IB})\tau+2\tilde{\psi}\right],$ (87g)
$c_{12}=$ $\vartheta_{61}\cos\left[2(\Omega+\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{62}\cos\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\vartheta_{63}\cos\left[2(\Omega-\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{64}\cos(2\omega_{IB}\tau)+\vartheta_{65},$ (87h)
$c_{12}^{\prime}=$ $\vartheta_{61}^{\prime}\sin\left[2(\Omega+\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{62}^{\prime}\sin\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\vartheta_{63}^{\prime}\sin\left[2(\Omega-\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{64}^{\prime}\sin(2\omega_{IB}\tau),$ (87i)
$c_{22}=$ $\vartheta_{71}\sin\left[(\Omega+2\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{72}\sin\left[(\Omega-2\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{73}\sin\left(\Omega\tau+\tilde{\psi}\right),$ (87j)
$c_{22}^{\prime}=$ $\vartheta_{71}^{\prime}\cos\left[(\Omega+2\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{72}^{\prime}\cos\left[(\Omega-2\omega_{IB})\tau+\tilde{\psi}\right]$
$+\vartheta_{73}^{\prime}\cos\left(\Omega\tau+\tilde{\psi}\right),$ (87k)
$c_{32}=$ $\vartheta_{81}\sin\left[2(\Omega+\omega_{IB})\tau+2\tilde{\psi}\right]$
$+\vartheta_{82}\sin\left(2\Omega\tau+2\tilde{\psi}\right)$
$+\vartheta_{83}\sin\left[2(\Omega-\omega_{IB})\tau+2\tilde{\psi}\right].$ (87l)
@@
$\vartheta_{11}=\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\left(1-\sin\tilde{\theta}\right),$ (88a)
$\vartheta_{12}=\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\left(1+\sin\tilde{\theta}\right),$ (88b)
$\vartheta_{21}=\frac{1}{4\gamma_{rr}^{1/2}}\left(1-\sin\tilde{\theta}\right),$ (88c)
$\vartheta_{22}=\frac{1}{4\gamma_{rr}^{1/2}}\left(1+\sin\tilde{\theta}\right),$ (88d)
$\vartheta_{21}^{\prime}=-(\Omega+\omega_{IB})\vartheta_{21},$ (88e)
$\vartheta_{22}^{\prime}=-(\Omega-\omega_{IB})\vartheta_{22},$ (88f)
$\vartheta_{31}=\frac{1}{2\gamma_{\theta\theta}^{1/2}}\cos\tilde{\theta},$ (88g)
$\vartheta_{31}^{\prime}=\omega_{IB}\vartheta_{31},$ (88h)
$\vartheta_{41}=\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\left(1-\sin\tilde{\theta}\right),$ (88i)
$\vartheta_{42}=\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\left(1+\sin\tilde{\theta}\right),$ (88j)
$\vartheta_{51}=\Theta_{0}\left(1-\sin\tilde{\theta}\right)^{2},$ (88k)
$\vartheta_{52}=\Theta_{0}\cos^{2}\tilde{\theta},$ (88l)
$\vartheta_{53}=\Theta_{0}\left(1+\sin\tilde{\theta}\right)^{2},$ (88m)
$\Theta_{0}=\frac{\left(\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu\right)\kappa_{\varphi}-\left(\gamma_{\varphi\varphi}^{1/2}+\mathcal{M}\mathcal{M}_{\varphi}\nu\right)\Omega}{32\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma},$ (88n)
$\vartheta_{61}=\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\left(1-\sin\tilde{\theta}\right)^{2},$ (88o)
$\vartheta_{62}=\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta},$ (88p)
$\vartheta_{63}=\frac{\kappa_{r}+\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\left(1+\sin\tilde{\theta}\right)^{2},$ (88q)
$\vartheta_{64}=\frac{\kappa_{r}+4\kappa_{\theta}-\kappa_{\varphi}}{64\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta},$ (88r)
$\vartheta_{65}=\frac{\kappa_{r}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\left(1+\sin^{2}\tilde{\theta}\right)-\frac{\kappa_{\theta}}{16\gamma_{rr}^{1/2}}\cos^{2}\tilde{\theta},$ (88s)
$\vartheta_{61}^{\prime}=-2(\Omega+\omega_{IB})\vartheta_{61},$ (88t)
@@
$\vartheta_{63}^{\prime}=-2(\Omega-\omega_{IB})\vartheta_{63},$ (88v)
$\vartheta_{64}^{\prime}=-2\omega_{IB}\vartheta_{64},$ (88w)
$\vartheta_{71}=\frac{\kappa_{\theta}\cos\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\left(1-\sin\tilde{\theta}\right),$ (88x)
$\vartheta_{72}=-\frac{\kappa_{\theta}\cos\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\left(1+\sin\tilde{\theta}\right),$ (88y)
$\vartheta_{73}=\frac{\kappa_{\theta}}{16\gamma_{\theta\theta}^{1/2}}\sin\left(2\tilde{\theta}\right),$ (88z)
$\vartheta_{71}^{\prime}=\left(\Omega+2\omega_{IB}\right)\vartheta_{71},$ (88aa)
$\vartheta_{72}^{\prime}=\left(\Omega-2\omega_{IB}\right)\vartheta_{72},$ (88ab)
$\vartheta_{73}^{\prime}=\Omega\vartheta_{73},$ (88ac)
$\vartheta_{81}=\frac{\kappa_{\varphi}-\nu\Omega}{32\gamma_{\varphi\varphi}^{1/2}/\gamma}\left(1-\sin\tilde{\theta}\right)^{2},$ (88ad)
$\vartheta_{82}=\frac{\kappa_{\varphi}-\nu\Omega}{32\gamma_{\varphi\varphi}^{1/2}/\gamma}\cos^{2}\tilde{\theta},$ (88ae)
$\vartheta_{83}=\frac{\kappa_{\varphi}-\nu\Omega}{32\gamma_{\varphi\varphi}^{1/2}/\gamma}\left(1+\sin\tilde{\theta}\right)^{2}.$ (88af)

### v2

@@ Appendix B Teukolsky source term
The source term in Eq. (12) is given by $\hat{T}=2\rho^{-4}T_{4}$, where $T_{4}$ is a combination of the projected stress-energy tensor defined by
@@ Appendix C Analytical expressions of split amplitudes
$Z^{(1)}_{lm}(\tau,\omega)=$$\big{\{}\tilde{\alpha}_{1}\sin\left[\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$ (82)
$+\tilde{\alpha}_{2}\cos\left[\left(\Omega+\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$+\tilde{\alpha}_{3}\sin\left[\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$+\tilde{\alpha}_{4}\cos\left[\left(\Omega-\omega_{\mathrm{IB}}\right)\tau-\tilde{\varphi}\right]$
$+\tilde{\alpha}_{5}\sin\left(\omega_{\mathrm{IB}}\tau\right)+\tilde{\alpha}_{6}\cos\left(\omega_{\mathrm{IB}}\tau\right)\big{\}}$
@@
$Z^{(1)}_{lm}(\tau,\omega)=$ $Z^{(1)}_{lm-1-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\Omega+\omega_{\mathrm{IB}}\right)\tau\right]$ (83)
$+Z^{(1)}_{lm11}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\Omega-\omega_{\mathrm{IB}}\right)\tau\right]$
$+Z^{(1)}_{lm-11}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\Omega-\omega_{\mathrm{IB}}\right)\tau\right]$
$+Z^{(1)}_{lm1-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\Omega+\omega_{\mathrm{IB}}\right)\tau\right]$
$+Z^{(1)}_{lm0-1}\exp\left[i\left(\omega u^{t}-mu^{\varphi}+\omega_{\mathrm{IB}}\right)\tau\right]$
$+Z^{(1)}_{lm01}\exp\left[i\left(\omega u^{t}-mu^{\varphi}-\omega_{\mathrm{IB}}\right)\tau\right].$
@@
$Z^{(1)}_{lm-1-1}=\frac{-i\tilde{\alpha}_{1}+\tilde{\alpha}_{2}}{2}e^{-i\tilde{\varphi}},$ (84a)
$Z^{(1)}_{lm-11}=\frac{-i\tilde{\alpha}_{3}+\tilde{\alpha}_{4}}{2}e^{-i\tilde{\varphi}},$ (84b)
@@
$Z^{(1)}_{lm1-1}=\frac{i\tilde{\alpha}_{3}+\tilde{\alpha}_{4}}{2}e^{+i\tilde{\varphi}},$ (84e)
$Z^{(1)}_{lm11}=\frac{i\tilde{\alpha}_{1}+\tilde{\alpha}_{2}}{2}e^{+i\tilde{\varphi}},$ (84f)
@@
$\tilde{\alpha}_{1}=$ $\frac{\gamma\left(1+\cos\tilde{\theta}\right)}{4\gamma_{\varphi\varphi}^{1/2}}\left[\alpha_{1}\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{\mathcal{M}}+\alpha_{4}\right]$
$-\frac{(\Omega+\omega_{\mathrm{IB}})(1+\cos\tilde{\theta})}{4\gamma_{rr}^{1/2}}\alpha_{5},$ (85a)
$\tilde{\alpha}_{2}=$ $\frac{1+\cos\tilde{\theta}}{4\gamma_{rr}^{1/2}}\alpha_{2},$ (85b)
$\tilde{\alpha}_{3}=$ $\frac{\gamma\left(1-\cos\tilde{\theta}\right)}{4\gamma_{\varphi\varphi}^{1/2}}\left[\alpha_{1}\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{\mathcal{M}}+\alpha_{4}\right]$
$-\frac{(\Omega-\omega_{\mathrm{IB}})(1-\cos\tilde{\theta})}{4\gamma_{rr}^{1/2}}\alpha_{5},$ (85c)
$\tilde{\alpha}_{4}=$ $\frac{1-\cos\tilde{\theta}}{4\gamma_{rr}^{1/2}}\alpha_{2},$ (85d)
$\tilde{\alpha}_{5}=$ $\frac{\sin\tilde{\theta}}{2\gamma_{\theta\theta}^{1/2}}\alpha_{3},$ (85e)
$\tilde{\alpha}_{6}=$ $\frac{\omega_{\mathrm{IB}}\sin\tilde{\theta}}{2\gamma_{\theta\theta}^{1/2}}\alpha_{6}.$ (85f)
@@
$c_{01}=$ $\vartheta_{11}\sin\left[(\Omega+\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{12}\sin\left[(\Omega-\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right],$ (87a)
$c_{11}=$ $\vartheta_{21}\cos\left[(\Omega+\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{22}\cos\left[(\Omega-\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right],$ (87b)
$c_{11}^{\prime}=$ $\vartheta_{21}^{\prime}\sin\left[(\Omega+\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{22}^{\prime}\sin\left[(\Omega-\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right],$ (87c)
$c_{21}=$ $\vartheta_{31}\sin(\omega_{\mathrm{IB}}\tau),$ (87d)
$c_{21}^{\prime}=$ $\vartheta_{31}^{\prime}\cos(\omega_{\mathrm{IB}}\tau),$ (87e)
$c_{31}=$ $\vartheta_{41}\sin\left[(\Omega+\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{42}\sin\left[(\Omega-\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right],$ (87f)
$c_{02}=$ $\vartheta_{51}\sin\left[2(\Omega+\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{52}\sin\left(2\Omega\tau-2\tilde{\varphi}\right)$
$+\vartheta_{53}\sin\left[2(\Omega-\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right],$ (87g)
$c_{12}=$ $\vartheta_{61}\cos\left[2(\Omega+\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{62}\cos\left(2\Omega\tau-2\tilde{\varphi}\right)$
$+\vartheta_{63}\cos\left[2(\Omega-\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{64}\cos(2\omega_{\mathrm{IB}}\tau)+\vartheta_{65},$ (87h)
$c_{12}^{\prime}=$ $\vartheta_{61}^{\prime}\sin\left[2(\Omega+\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{62}^{\prime}\sin\left(2\Omega\tau-2\tilde{\varphi}\right)$
$+\vartheta_{63}^{\prime}\sin\left[2(\Omega-\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{64}^{\prime}\sin(2\omega_{\mathrm{IB}}\tau),$ (87i)
$c_{22}=$ $\vartheta_{71}\cos\left[(\Omega+2\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{72}\cos\left[(\Omega-2\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{73}\cos\left(\Omega\tau-\tilde{\varphi}\right),$ (87j)
$c_{22}^{\prime}=$ $\vartheta_{71}^{\prime}\sin\left[(\Omega+2\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{72}^{\prime}\sin\left[(\Omega-2\omega_{\mathrm{IB}})\tau-\tilde{\varphi}\right]$
$+\vartheta_{73}^{\prime}\sin\left(\Omega\tau-\tilde{\varphi}\right),$ (87k)
$c_{32}=$ $\vartheta_{81}\sin\left[2(\Omega+\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right]$
$+\vartheta_{82}\sin\left(2\Omega\tau-2\tilde{\varphi}\right)$
$+\vartheta_{83}\sin\left[2(\Omega-\omega_{\mathrm{IB}})\tau-2\tilde{\varphi}\right].$ (87l)
@@
$\vartheta_{11}=\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\left(1+\cos\tilde{\theta}\right),$ (88a)
$\vartheta_{12}=-\frac{\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu}{4\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma}\left(1-\cos\tilde{\theta}\right),$ (88b)
$\vartheta_{21}=\frac{1}{4\gamma_{rr}^{1/2}}\left(1+\cos\tilde{\theta}\right),$ (88c)
$\vartheta_{22}=-\frac{1}{4\gamma_{rr}^{1/2}}\left(1-\cos\tilde{\theta}\right),$ (88d)
$\vartheta_{21}^{\prime}=-(\Omega+\omega_{\mathrm{IB}})\vartheta_{21},$ (88e)
$\vartheta_{22}^{\prime}=-(\Omega-\omega_{\mathrm{IB}})\vartheta_{22},$ (88f)
$\vartheta_{31}=\frac{1}{2\gamma_{\theta\theta}^{1/2}}\sin\tilde{\theta},$ (88g)
$\vartheta_{31}^{\prime}=-\omega_{\mathrm{IB}}\vartheta_{31},$ (88h)
$\vartheta_{41}=\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\left(1+\cos\tilde{\theta}\right),$ (88i)
$\vartheta_{42}=-\frac{\gamma}{4\gamma_{\varphi\varphi}^{1/2}}\left(1-\cos\tilde{\theta}\right),$ (88j)
$\vartheta_{51}=2\Theta_{0}\cos^{4}\frac{\tilde{\theta}}{2},$ (88k)
$\vartheta_{52}=-\Theta_{0}\sin^{2}\tilde{\theta},$ (88l)
$\vartheta_{53}=2\Theta_{0}\sin^{4}\frac{\tilde{\theta}}{2},$ (88m)
$\Theta_{0}=\frac{\left(\mathcal{M}\mathcal{M}_{\varphi}+\gamma_{\varphi\varphi}^{1/2}\nu\right)\kappa_{\varphi}-\left(\gamma_{\varphi\varphi}^{1/2}+\mathcal{M}\mathcal{M}_{\varphi}\nu\right)\Omega}{16\gamma_{\varphi\varphi}^{1/2}\mathcal{M}/\gamma},$ (88n)
$\vartheta_{61}=\frac{\kappa_{r}+\kappa_{\varphi}}{16\gamma_{rr}^{1/2}}\cos^{4}\frac{\tilde{\theta}}{2},$ (88o)
$\vartheta_{62}=\frac{\kappa_{r}+\kappa_{\varphi}}{16\gamma_{rr}^{1/2}}\sin^{4}\frac{\tilde{\theta}}{2},$ (88p)
$\vartheta_{63}=-\frac{\kappa_{r}+\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta},$ (88q)
$\vartheta_{64}=-\frac{\kappa_{r}+2\kappa_{\theta}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta},$ (88r)
$\vartheta_{65}=\frac{\kappa_{r}-\kappa_{\varphi}}{32\gamma_{rr}^{1/2}}\left(1+\cos^{2}\tilde{\theta}\right)-\frac{\kappa_{\theta}}{16\gamma_{rr}^{1/2}}\sin^{2}\tilde{\theta},$ (88s)
$\vartheta_{61}^{\prime}=-2(\Omega+\omega_{\mathrm{IB}})\vartheta_{61},$ (88t)
@@
$\vartheta_{63}^{\prime}=-2(\Omega-\omega_{\mathrm{IB}})\vartheta_{63},$ (88v)
$\vartheta_{64}^{\prime}=-2\omega_{\mathrm{IB}}\vartheta_{64},$ (88w)
$\vartheta_{71}=\frac{\kappa_{\theta}\cos\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\left(1+\cos\tilde{\theta}\right),$ (88x)
$\vartheta_{72}=-\frac{\kappa_{\theta}\cos\tilde{\theta}}{16\gamma_{\theta\theta}^{1/2}}\left(1-\cos\tilde{\theta}\right),$ (88y)
$\vartheta_{73}=\frac{\kappa_{\theta}}{8\gamma_{\theta\theta}^{1/2}}\cos\left(2\tilde{\theta}\right),$ (88z)
$\vartheta_{71}^{\prime}=-\left(\Omega+2\omega_{\mathrm{IB}}\right)\vartheta_{71},$ (88aa)
$\vartheta_{72}^{\prime}=-\left(\Omega-2\omega_{\mathrm{IB}}\right)\vartheta_{72},$ (88ab)
$\vartheta_{73}^{\prime}=-\Omega\vartheta_{73},$ (88ac)
$\vartheta_{81}=\frac{\kappa_{\varphi}-\nu\Omega}{8\gamma_{\varphi\varphi}^{1/2}/\gamma}\cos^{4}\frac{\tilde{\theta}}{2},$ (88ad)
$\vartheta_{82}=-\frac{\kappa_{\varphi}-\nu\Omega}{16\gamma_{\varphi\varphi}^{1/2}/\gamma}\sin^{2}\tilde{\theta},$ (88ae)
$\vartheta_{83}=\frac{\kappa_{\varphi}-\nu\Omega}{6\gamma_{\varphi\varphi}^{1/2}/\gamma}\sin^{4}\frac{\tilde{\theta}}{2}.$ (88af)